Dear Sir, For the function f(x)=X^x . Find the points at which it is 1.Increasing. 2. Decreasing . Hence Determine which of e^(pi) , (pi)^ e is greater .

Dear Sir,


For the function f(x)=X^x . Find the points at which it is


1.Increasing.


2. Decreasing .


Hence Determine which of e^(pi)  , (pi)^ e  is greater .

Grade:11

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Increasing Functions

A function is "increasing" if the y-value increases as the x-value increases, like this:

Increasing Function

It is easy to see that y=f(x) tends to go up as it goes along.

What about that flat bit near the start? Is that OK?

  • Yes, it is OK if you say the function is Increasing
  • But it is not OK if you say the function is Strictly Increasing (no flatness allowed)

Using Algebra

What if you can't plot the graph to see if it is increasing? In that case is is good to have a definition using algebra.

For a function y=f(x):

when x1 < x2 then f(x1) ≤ f(x2) Increasing
when x1 < x2 then f(x1) < f(x2) Strictly Increasing

That has to be true for any x1, x2, not just some nice ones you choose.

 

vikas askiitian expert
13 years ago

f(x) = xx

 

d/dx [f(x)] = d/dx [ xx ]

 

 f1(x)       =xx [ 1 + logx ]

f1(x) = 0 , when 1+logx = 0 ,  xx cannot be 0..

    1+ logx = 0

      logx = -1

        x = 1/e 

if x > 1/e , then function is increasing

if x < 1/e , then function is decreasing

increasing => (1/e , infinity)

decreasing=> (-infinity,1/e)

jagdish singh singh
13 years ago